OpenStudy (samsan9):

a baseball is thrown directly upward with an initial speed of 25 m/s will rise to what height?

OpenStudy (samsan9):

@stefrheart

OpenStudy (stefrheart):

okay

OpenStudy (anonymous):

use this formula,\[v^2 = u^2 - 2gh \]

OpenStudy (samsan9):

ok so um what do i plug in because my teacher hasnt really covered that

OpenStudy (anonymous):

well yeah plug in the value as v=0 and rest of all is given like, u=25m/s g=9.8 and "h" u have to calculate....

OpenStudy (samsan9):

490?

OpenStudy (anonymous):

nopes...that's not the answer....

OpenStudy (anonymous):

0=625-9.8*2*h h=625/19.6= 31.8( approx 32 m)

OpenStudy (samsan9):

thank you :)

OpenStudy (anonymous):

np ) anytime :D @samsan9